OpenStudy (anonymous):

Compare and contrast these four functions using complete sentences. Focus on the slope and y-intercept first, and then on any additional properties of each function.

OpenStudy (anonymous):

click the link to see a picture of the functions

OpenStudy (zpupster):

Live preview of your reply: lets take the first function f(x)-- he gets 5$ from his dad that is a constant--+5 next he gets 2$ every time walking the dog--+2 how many times does he walk the dog--X so f(x) = 2x+5

OpenStudy (zpupster):

now look at j(x) and compare with f(x)

OpenStudy (zpupster):

now H(x) from the graph we see that when x is 0 y or h(x) is 5 so our y intercept is 5 so y = ?x +5 now when y is 0 , x is about 2 1/2 or 5/2 solving 0 = 2x + 5 0-5 = 2x + 5 -5 subtracting 5 both sides -5/2 = 2x/2 dividing 2 both sides we get x= -5/2 but we need positive 5/2 so the 2 has to be negative h(x) = -2x+5 compare this with the other 2 functions

OpenStudy (zpupster):

i will let you do g(x), just take the function and plug in the x value and see which function will satisfy the equation --equalling g(x) also you can put your compare and contrast into words.

OpenStudy (anonymous):

@zpupster so g(x)=-1x-2??

OpenStudy (zpupster):

you have to find an equation that so when you place the x it satisfy to g(x) or y so y=mx+b the x is -2 and y or g(x) is -1 -1 = -2(-2) -5 so now y or g(x) = -9 -9 = -2(2) -5 and lastly when y = -17 when x is 6 -17 = -2(6) -5 now you see the equation for g(x)

OpenStudy (anonymous):

ooooh okay ! @zpupster

OpenStudy (anonymous):

thank you !